The following costs were for Optimal View Inc., a contact lens manufacturer:

Output Fixed Cost Variable Cost Total Costs
270 $ 5,300 $ 10,000 $ 15,300
320 5,300 12,000 17,300
370 5,300 14,000 19,300
420 5,300 16,000 21,300

At an output level of 500 lenses, per unit total cost is projected to be:

Multiple Choice

  • $44.45.

  • $49.00.

  • $38.05.

  • $55.27.

  • $47.16.
Answer #1

Variable cost per unit = 10,000/270 = 37.03

Total Cost for 500 units = (500*37.03) + 5,300 = 23, 845

Cost per unit = 23,845/500 = 47.69

OPTION E IS THE ANSWER (nearest number)